Find the work done in stretching a spring of natural length 8 cm, from 10 cm to 13 cm. Assume a force of 6 N is needed to hold it at a length of 11 cm.
1
Expert's answer
2020-04-06T16:13:15-0400

"Force=Spring\ constant\ *\ elongation\\
F=k*e\\"

"at\ length=11cm,\\
6=k*(11-8)\\
k=2Ncm^{-1}\\"

"at\ length=10cm,\\
F_{10cm}=2*(10-8)\\
F_{10cm}=4N\\"

"at\ length=13cm,\\
F_{13cm}=2*(13-8)\\
F_{13cm}=10N\\"

"work\ done\ =\ force\ *\ distance\\
work\ done\ =\ \frac{(4+10)}{2}\ *\ (13-10)*10^{-2}\\ (multiply\ by\ 10^{-2}\ for\ convert\ into\ meters.)\\
\bold {work\ done\ =\ 0.21Nm}"